# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

### v2

@@ 4.3 Heterophily feature weighting
In recommender systems, users and items exhibit diverse characteristics. For instance, in movie recommender systems, individual movies, as items, possess unique attributes. Movies vary in terms of cast members, with differing numbers of renowned actors featured. Additionally, movies belong to distinct genres, and some may even encompass multiple genres. Consequently, we exploit homophily rate to understand users’ preferences for these diverse items’ characteristics. To achieve this goal, we construct an item-item graph, linking items within a users’ profile. Each item is assigned its category as a label. Table 1 summarizes the item-item graph data statistics [^1] In Section 5, we use six datasets for our empirical evaluations. However, in Table 1, only five datasets are reported. The reason is that the Epinions dataset lacks the category information necessary for constructing an item-item graph. Therefore, we do not report it in Table 1.. The homophily rate is determined using the GEOM-GCN method [22].
Table 1: Item-item graph statistics.
Dataset #nodes #edges #classes Homophily
Amazon-Electronics 11589 3790708 528 0.03
Amazon-Beauty 733 119329 112 0.05
Amazon-Health 1260 379736 290 0.01
Amazon-Movie 21901 43758979 77 0.7
Amazon-CDs 16184 18113968 209 0.11
We notice from Table 1 that:
  • Low homophily rate indicates that the nodes in the graph have diverse attributes and characteristics. In essence, there is a limited tendency for nodes to connect with others sharing similar features. This observation indicates a high level of diversity within the network.
  • Nodes may form connections based on a variety of factors, and the network might be characterized by a mixture of different classes or categories. This diversity could be beneficial for information diffusion or collaboration across various domains.
Based on the above analysis, instead of assigning a unique weight $\sigma(h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)})$ to the message between $u$ and $i$, we opt for distinct weights for different features in each propagated message. This enables us to effectively capture dissimilarities between user and item features. So, we define the hererophily feature weighting as follows:
@@
where $W^{\left(k\right)}\in\mathbb{R}^{1\times d}$ is a trainable linear transformation, $d$ is the dimension of node embeddings and $w_{ui}^{(k)}\in\mathbb{R}^{1\times d}$. Therefore, Equations 11 and 12 are revised as follows:
@@ 5.1.1 Datasets
We examine the algorithms over six well-known datasets: Epinions [^2] https://snap.stanford.edu/data/soc-Epinions1.html [42], Amazon-Movies [^3] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html [13], Amazon-Electronics3 [13], Amazon-Health3 [13], Amazon-Beauty3 [13] and Amazon-CDs3 [13]. For the amazon review datasets, we adopt a $10$-core setting, filtering out users and items with less than $10$ interactions. In the case of Epinions, we utilize the data released by APDA  [42]. Table 2 depicts the statistics of the datasets.
@@ 5.1.2 Baselines
  • PC [44]: This is a pos-processing (ReRanking) approach that manipulates predicted scores to calculate new scores as a function of item popularity. We tune the parameter $\alpha$ and $\beta$ as described in the original paper.
@@ 5.1.3 Evaluation metrics
  • Statistical Parity (or Demographic Parity): this approach necessitates that different groups of items or similar individuals are treated similarly, without considering users’ interests. This approach fails to consider the users’ side and increases the risk of damaging users’ satisfaction.
  • Equal Opportunity: this approach mandates that groups of items benefiting from the outcome are afforded equal opportunities. It strives to treat all items fairly within users’ profiles, regardless of their popularity.
Researchers in [44] introduce two metrics based on equal opportunity:
@@
Here, $SRC\left(\cdot,\cdot\right)$ calculates Spearman’s rank correlation, $\tilde{O}_{u}^{+}$ represents items in the test profile of user $u$, $rank_{u}\left(x\right)$ signifies the ranking of item $x$ that the model predicts for user $u$, $U_{i}$ is a set of users that item $i$ already exists in their test profiles and $d_{x}$ is the degree of node $x$. While PRU investigates items’ position bias for individual users’ recommendation lists, PRI considers the items’ side fairness across all users. A large value of the latter may result in long-tail items having limited opportunities to become popular [44]. In general, the methods’ performance in promoting fairness improves as the values of these two criteria get closer to 00.
@@ 5.1.4 Hyper-parameters setting
For all methods, we use the same early stopping mechanism, stopping training after $15$ epochs if NGCF does not show improvement in these $15$ epochs. The embeddings dimension is set to $64$ for LightGCN and AdjNorm, $256$ for the APDA model, and $128$ for our model. The only exception is the Movies dataset where due to hardware limitations, we set the embeddings dimension to 64. The batch size for LightGCN and AdjNorm is set to $1024$, as proposed in their papers, while for APDA we set it to $2048$. The learning rate for all the baselines is set to $0.001$. For all the methods, we set both the regularization hyper-parameter $\beta$ and the weight decay parameter to $0.0001$. The learning rate for our model is set to $0.0005$. We set the number of propagation layers to $3$ in LightGCN and r-AdjNorm, and to $4$ in APDA and HetroFair.
@@ 5.2 Performance comparison
Table 3 shows the results of HetroFair in comparison to the other methods. The best performance is marked in bold, and the second-best performance is underlined. The findings indicate that our method outperforms the other methods, not only in the fairness metrics but also in the accuracy metrics. These results show that our method is better fitted with the data and takes into account their distinct features. In the case of baselines, it is observable that although the Reg method often achieves commendable fairness metrics compared to other baselines, its accuracy is significantly lower than that of other methods. This suggests that regularization-based methods may reduce the performance of recommendations for end-users.
@@
Dataset Metric LightGCN LightGCN+Reg LightGCN+PC r-AdjNorm APDA Our model w/o Hetro
Epinions NDCG 0.0789 0.0746 0.0778 0.0798 0.0859 0.0895 0.0889
MRR 0.1361 0.1288 0.1336 0.1369 0.1471 0.1525 0.1522
MAP 0.0329 0.0307 0.0324 0.0339 0.0366 0.0379 0.0377
PRU 0.5233 0.5114 0.5168 0.5077 0.5188 0.4706 0.5054
PRI 0.5129 0.5065 0.5071 0.5366 0.4983 0.4384 0.4573
Electronics NDCG 0.0463 0.0407 0.0422 0.0461 0.0517 0.0525 0.0488
MRR 0.0645 0.0552 0.0610 0.0640 0.0708 0.0733 0.0676
MAP 0.0221 0.0188 0.0203 0.0219 0.0250 0.0256 0.0232
PRU 0.5216 0.4779 0.4919 0.5300 0.5548 0.4600 0.5020
PRI 0.5417 0.5077 0.5325 0.5661 0.5692 0.4462 0.4805
CDs NDCG 0.1231 0.1230 0.1245 0.1178 0.1379 0.1449 0.1346
MRR 0.1715 0.1723 0.1760 0.1638 0.1940 0.2017 0.1882
MAP 0.0611 0.0611 0.0617 0.0563 0.0710 0.0747 0.0680
PRU 0.3304 0.3287 0.3270 0.3262 0.3265 0.2972 0.3266
PRI 0.2824 0.3069 0.2974 0.3193 0.2763 0.2264 0.2495
Health NDCG 0.1259 0.1202 0.1224 0.1271 0.1314 0.1334 0.1351
MRR 0.1986 0.1902 0.196 0.2028 0.2073 0.2112 0.2093
MAP 0.0595 0.0570 0.0574 0.0610 0.0637 0.0656 0.0669
PRU 0.5162 0.4980 0.4893 0.5270 0.5205 0.4922 0.4830
PRI 0.5303 0.5116 0.5285 0.5240 0.5555 0.5028 0.4804
Beauty NDCG 0.2113 0.2161 0.2116 0.2136 0.2274 0.2308 0.2302
MRR 0.2525 0.2603 0.2540 0.2520 0.2790 0.2824 0.2798
MAP 0.1200 0.1240 0.1203 0.1235 0.1354 0.1364 0.1389
PRU 0.4428 0.4407 0.04385 0.4188 0.4418 0.4139 0.3117
PRI 0.2164 0.2071 0.2151 0.2089 0.2266 0.1821 0.1270
Movies NDCG 0.0672 0.0673 0.0690 0.0676 0.0715 0.0777 0.0741
MRR 0.0966 0.0957 0.0971 0.0955 0.1011 0.1093 0.1048
MAP 0.0304 0.0298 0.0308 0.0299 0.0331 0.0365 0.0340
PRU 0.3098 0.3051 0.3082 0.3245 0.3384 0.3045 0.3251
PRI 0.3951 0.3714 0.3829 0.4436 0.3862 0.3355 0.3590
@@ 5.4 Hyper-parameters study
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 2 and 3. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN and PC models drastically declines after 5 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. Considering the NDCG metric in Figures 4 and 5, the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity.
@@
Parameter $\lambda$ plays a crucial role in the accuracy-fairness trade-off. We consider the datasets Health and Beauty, and change the value of $\lambda$ from $0.05$ to $0.95$ with $0.05$ step. The results are depicted in Figures 6 and 7. Over both datasets, we observe almost the same pattern: within first few steps, the fairness metrics are in their lowest values, while the accuracy metrics are also too low. As the value of $\lambda$ increases, the performance of HetroFair increases, while PRU and RRI deteriorate until reaching to an optimal value for $\lambda$, which leads to almost stable values for all the metrics.
@@ 6 Conclusion
In this paper, we studied items’ side fairness in GNN-based recommendation systems. First, we pointed out that message normalization and aggregation processes in GNNs could lead to popularity bias. Then, we proposed the HetroFair model, that utilizes two techniques to overcome this bias and generate fair embeddings. The first technique was fairness-aware attention, which incorporates dot product in the normalization process of GNNs, to decrease the effect of nodes’ degrees. The second technique was heterophily feature weighting, which assigns distinct weights to different features during the aggregation process. Our extensive experiments over six well-known datasets demonstrated the high performance of HetroFair, in improving both users’ side accuracy and items’ side fairness.

### v3

@@ 4.3 Heterophily feature weighting
The similarity function employed in the previous section is the dot product, a widely used measure for quantifying the alignment between two vectors. Despite its simplicity and computational efficiency, the dot product has a notable limitation: it is not trainable. This lack of trainability means it cannot adapt to the characteristics of the training data or optimize itself for specific tasks during the learning process. Consequently, it may fail to capture subtle relationships between users and items, which are critical for improving performance in recommendation tasks. To overcome this challenge, we propose a straightforward yet effective modification to the dot product, transforming it into a trainable similarity function. This enhancement enables the model to dynamically adjust similarity computations based on the training data.
In our approach, we introduce feature-specific weights to the user and item vectors, denoting them as $w_{1}$ for the first feature in users and items vector, $w_{2}$ for second feature and so forth. These weights are not predefined but are learned adaptively during the training process. This process is integrated into the message-passing layers, allowing the model to incorporate feature-specific adjustments at every stage of information propagation. We call this method the heterophily-aware similarity function because it operates at the feature level and assigns distinct, non-equal weights to capture heterophilous relationships. So, instead of assigning a unique weight $\sigma(h_{u}^{\left(k\right)}\cdot h_{i}^{\left(k\right)})$ to the message between $u$ and $i$, we define the heterophily-aware feature weighting as follows:
@@
where $W^{\left(k\right)}\in\mathbb{R}^{1\times d}$ is a trainable linear transformation, $d$ is the dimension of node embeddings and $w_{ui}^{(k)}\in\mathbb{R}^{1\times d}$. Therefore, Equations 8 and 9 are revised as follows:
@@ 5.1.1 Datasets
We examine the algorithms over six well-known datasets: Epinions [^1] https://snap.stanford.edu/data/soc-Epinions1.html [42], Amazon-Movies [^2] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html [13], Amazon-Electronics2 [13], Amazon-Health2 [13], Amazon-Beauty2 [13] and Amazon-CDs2 [13]. For the amazon review datasets, we adopt a $10$-core setting, filtering out users and items with less than $10$ interactions. In the case of Epinions, we utilize the data released by APDA  [42]. Table 1 depicts the statistics of the datasets.
@@ 5.1.2 Baselines
  • PC [44]: This is a pos-processing (re-ranking) approach that manipulates predicted scores to calculate new scores as a function of item popularity. We tune the parameter $\alpha$ and $\beta$ as described in the original paper.
@@
  • PopGo [37]: This method mitigates popularity bias by training a shortcut model to estimate interaction-wise shortcut degrees and adjusting CF model predictions accordingly. It quantifies shortcuts based on popularity representations of user-item pairs. Following the original paper, we fine-tune the hyperparameters $\tau_{1}$ and $\tau_{2}$ within the recommended ranges to ensure optimal performance.
@@ 5.1.3 Evaluation metrics
  • Statistical Parity (or Demographic Parity): this metric necessitates that different groups of items or similar individuals are treated similarly, without considering users’ interests. This approach fails to consider the users’ side and increases the risk of damaging users’ satisfaction.
  • Equal Opportunity: this metric mandates that groups of items benefiting from the outcome are afforded equal opportunities. It strives to treat all items fairly within users’ profiles, regardless of their popularity.
The authors of [44] introduce two metrics based on equal opportunity:
@@
Here, $SRC\left(\cdot,\cdot\right)$ calculates Spearman’s rank correlation, $\tilde{O}_{u}^{+}$ represents items in the test profile of user $u$, $rank_{u}\left(x\right)$ signifies the ranking of item $x$ that the model predicts for user $u$, $U_{i}$ is a set of users that item $i$ already exists in their test profiles and $d_{x}$ is the degree of node $x$. While PRU investigates items’ position bias for individual users’ recommendation lists, PRI considers item-side fairness across all users. A large value of the later may result in long-tail items having limited opportunities to become popular [44]. In general, the methods’ performance in promoting fairness improves as the values of these two criteria get closer to 00.
@@ 5.1.4 Hyper-parameters setting
For all methods, we use the same early stopping mechanism, stopping training after $15$ epochs if NGCF does not show improvement in these $15$ epochs. The embeddings dimension is set to $64$ for LightGCN and AdjNorm, $256$ for the APDA model, and $128$ for our model. The only exception is the Movies dataset where due to hardware limitations, we set the embeddings dimension to 64. The batch size for LightGCN and AdjNorm is set to $1024$, as proposed in their papers, while for APDA we set it to $2048$. The learning rate for all the baselines is set to $0.001$. For all the methods, we set both the regularization hyper-parameter $\beta$ and the weight decay parameter to $0.0001$. The learning rate for our model is set to $0.0005$. We set the number of propagation layers to $3$ in LightGCN and r-AdjNorm, and to $4$ in APDA and HetroFair. For the PopGo model, the optimal number of layers varies between $2$ and $3$ across different datasets. The embedding size is set to $64$. For the two hyperparameter $\tau_{1}$ and $\tau_{2}$, various values are tested within the same range specified in the paper to achieve the best results.
@@ 5.2 Performance comparison
Table 2 shows the results of HetroFair in comparison to the other methods. Over each dataset and for each metric, the best performance is marked in bold, and the second-best performance is underlined. The findings indicate that our method outperforms the other methods, not only in the fairness metrics but also in the accuracy metrics. Only in the two small and dense datasets (Health and Beauty), the PopGo model performs well on certain metrics. However, our method also performs well on these datasets while achieving a better balance between fairness and accuracy. These results show that our method is better fitted with the data and takes into account their distinct features. In the case of baselines, it is observable that although the Reg method often achieves commendable fairness metrics compared to other baselines, its accuracy is significantly lower than that of other methods. This suggests that regularization-based methods may reduce the performance of recommendations for end-users. Regarding the PopGo model, although it improves PRI in some cases, it fails to simultaneously enhance PRU, indicating its limited ability to adjust item rankings at the top of the recommendation list.
@@
Dataset Metric LightGCN LightGCN+Reg LightGCN+PC r-AdjNorm APDA PopGo Our model w/o Hetro
Epinions NDCG 0.0789 0.0746 0.0778 0.0798 0.0859 0.0879 0.0895 0.0889
MRR 0.1361 0.1288 0.1336 0.1369 0.1471 0.1515 0.1525 0.1522
MAP 0.0329 0.0307 0.0324 0.0339 0.0366 0.0376 0.0379 0.0377
PRU 0.5233 0.5114 0.5168 0.5077 0.5188 0.5976 0.4706 0.5054
PRI 0.5129 0.5065 0.5071 0.5366 0.4983 0.5475 0.4384 0.4573
Electronics NDCG 0.0463 0.0407 0.0422 0.0461 0.0517 0.0495 0.0525 0.0488
MRR 0.0645 0.0552 0.0610 0.0640 0.0708 0.0687 0.0733 0.0676
MAP 0.0221 0.0188 0.0203 0.0219 0.0250 0.0236 0.0256 0.0232
PRU 0.5216 0.4779 0.4919 0.5300 0.5548 0.5208 0.4600 0.5020
PRI 0.5417 0.5077 0.5325 0.5661 0.5692 0.5086 0.4462 0.4805
CDs NDCG 0.1231 0.1230 0.1245 0.1178 0.1379 0.1228 0.1449 0.1346
MRR 0.1715 0.1723 0.1760 0.1638 0.1940 0.1752 0.2017 0.1882
MAP 0.0611 0.0611 0.0617 0.0563 0.0710 0.0607 0.0747 0.0680
PRU 0.3304 0.3287 0.3270 0.3262 0.3265 0.4487 0.2972 0.3266
PRI 0.2824 0.3069 0.2974 0.3193 0.2763 0.2530 0.2264 0.2495
Health NDCG 0.1259 0.1202 0.1224 0.1271 0.1314 0.1311 0.1334 0.1351
MRR 0.1986 0.1902 0.196 0.2028 0.2073 0.2044 0.2112 0.2093
MAP 0.0595 0.0570 0.0574 0.0610 0.0637 0.0644 0.0656 0.0669
PRU 0.5162 0.4980 0.4893 0.5270 0.5205 0.4075 0.4922 0.4830
PRI 0.5303 0.5116 0.5285 0.5240 0.5555 0.3997 0.5028 0.4804
Beauty NDCG 0.2113 0.2161 0.2116 0.2136 0.2274 0.2382 0.2308 0.2302
MRR 0.2525 0.2603 0.2540 0.2520 0.2790 0.2870 0.2824 0.2798
MAP 0.1200 0.1240 0.1203 0.1235 0.1354 0.1415 0.1364 0.1389
PRU 0.4428 0.4407 0.04385 0.4188 0.4418 0.4091 0.4139 0.3117
PRI 0.2164 0.2071 0.2151 0.2089 0.2266 0.1900 0.1821 0.1270
Movies NDCG 0.0672 0.0673 0.0690 0.0676 0.0715 0.0682 0.0777 0.0741
MRR 0.0966 0.0957 0.0971 0.0955 0.1011 0.0979 0.1093 0.1048
MAP 0.0304 0.0298 0.0308 0.0299 0.0331 0.0314 0.0365 0.0340
PRU 0.3098 0.3051 0.3082 0.3245 0.3384 0.5034 0.3045 0.3251
PRI 0.3951 0.3714 0.3829 0.4436 0.3862 0.2818 0.3355 0.3590
@@ 5.4 Hyper-parameters study
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN and PC models drastically declines after 5 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 6 and 7. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. Considering the NDCG metric in Figures 6 and 7, the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity.
@@
Parameter $\lambda$ plays a crucial role in the accuracy-fairness trade-off. We consider the datasets Health and Beauty, and change the value of $\lambda$ from $0.05$ to $0.95$ with $0.05$ step. The results are depicted in Figures 8 and 9. Over both datasets, we observe almost the same pattern: within first few steps, the fairness metrics are in their lowest values, while the accuracy metrics are also too low. As the value of $\lambda$ increases, the performance of HetroFair increases, while PRU and RRI deteriorate until reaching to an optimal value for $\lambda$, which leads to almost stable values for all the metrics.
@@ 6 Conclusion
In this paper, we studied item-side fairness in GNN-based recommendation systems. First, we pointed out that message normalization and aggregation processes in GNNs could lead to popularity bias. Then, we proposed the HetroFair model, which utilizes two techniques to overcome this bias and generate fair embeddings. The first technique is fairness-aware attention, which incorporates the dot product in the normalization process of GNNs to decrease the effect of nodes’ degrees. The second technique is a trainable dot product similarity function, which assigns distinct weights to different features during the aggregation process. Our extensive experiments over six well-known datasets demonstrated the high performance of HetroFair in improving both user-side accuracy and item-side fairness.
Acknowledgment
This work is supported by the Iran National Science Foundation (INSF) under project No.4034377.